OpenStudy (anonymous):

What is the volume of grain that could completely fill this silo rounded to the nearest whole number? Use 22/7 for pi.

OpenStudy (anonymous):

13,750 ft3 14,012 ft3 262 ft3 4,583 ft3

Directrix (directrix):

Need a diagram or more info about the silo.

Directrix (directrix):

The silo is a right circular cylinder and a hemisphere. Volume of Cylinder =========== V = pi r^2 h V = pi (5)^2(175) V =4375 pi Volume of Hemisphere ============= Volume = (1/2) (4/3) pi (r)^3 Volume = (2/3) pi (5)^3 Volume = 250/3 pi Volume of Silo = 4458.33 pi = 14011.9 14,012 ft3 ------> ANSWER

OpenStudy (anonymous):

how did u get: Volume of Silo = 4458.33 pi = 14011.9